# ]The authors are members of LabEx Tec21 (Investissements d’Avenir - grant agreement ANR-11-LABX-0030)

### v2

@@ I Introduction
Red blood cell (RBC) aggregation is a critical factor influencing blood rheology and circulation. Due to the presence of plasma proteins, RBCs tend to reversibly aggregate, forming mostly linear structures known as rouleaux favored by their flat biconcave shape. These linear structures can assemble into larger, gel-like networks at low shear rates and are mostly responsible for the strong shear-thinning of blood observed at low shear rates chien1970 , in conjunction with RBC deformability chien1967a ; lanotte16 .
@@
There is no consensus on a strict equivalence between red cell density and age and studies show that there is a wide distribution of densities among cells of the same age, it is still accepted that on average density increases with cell age morrison1983does . However, a robust feature is that cell density is related to hemoglobin content mohandas86scattering , which has a direct and strong impact on cytoplasm viscosity. In addition, density increase was found to be correlated to loss of membrane surface area. Both phenomena lead to a decrease of static deformability (deformability index) and dynamic deformability (increase of relaxation time). Besides, density increase is also observed in RBC disorders associated with decreased cell deformability mohandas88densdistrib .
Recently, several studies have used Percoll or iodixanol density gradients to separate RBCs by density and analyze their dynamics, demonstrating a significant impact on RBC shape in flow nouaman2023effect , as well as migration and axial dispersion in capillaries losserand2019migration ; losserand2023axial .
Various techniques have been proposed to quantify aggregation and characterize the aggregability of erythrocytes at the suspension scale. These include the classic erythrocyte sedimentation rate (ESR) test bedell1985 , ektacytometry dacosta2016 , LORCA hardeman1994laser , and ultrasound scattering franceschini2010ultrasound , all of which are applied to dense suspensions. At the cellular scale, force measurements or controlled force application methods, such as micropipette aspiration techniques tozeren1989theoretical ; buxbaum1982quantitation , AFM steffen2013quantification , and more recently optical tweezers (OT) lee2017assessment ; Ermolinskiy2020Oct , allow direct investigation of cell-cell interactions. Notably, Ermolinskiy et al. Ermolinskiy2020Oct showed that the aggregation force measured by OT increases with the density of erythrocytes, both in autologous plasma and in 50 mg/ml Dextran 70 kDa solutions, suggesting an increase of RBC aggregability with cell age.
@@
In this work, we use this technique to quantify the variations in dissociation probabilities of small aggregates of healthy erythrocytes in extensional flow as a function of cell density after separating them using Percoll density gradients. We show that aggregates formed from the denser fractions of red blood cells are significantly more robust than the average cell population and require significantly higher flow stresses for dissociation. These results suggest that in addition to decreased deformability, RBC aging leads to increased aggregation in microcirculatory conditions.
@@ II.1 Blood sample collection and preparation
The separation of RBCs according to cell age was conducted by utilizing the density variations that arise during the ageing process Waugh1992Mar . To obtain density fractions of RBCs from a whole blood sample, we employed the Percoll density gradient centrifugation method outlined by Ermolinskiy et al. Ermolinskiy2020Oct . The Percoll solution (Cytiva 17-0891-01, Sigma-Aldrich, Taufkirchen, Germany) consists of Percoll, distilled water, and a 1.5M NaCl solution in different proportions, which creates the five different density levels for the experiment. Five solutions, with densities of 1.085 g/mL, 1.092 g/mL, 1.101 g/mL, 1.107 g/mL, and 1.122 g/mL, were prepared and carefully placed one over the other in a 15 mL centrifugation tube (highest density solution at the bottom, lowest at the top, with each layer of volume of 2 mL). Finally, a 2 mL suspension of washed red blood cells, diluted to a hematocrit of 60% in PBS, was carefully layered onto the Percoll gradient solution. The sample was then centrifuged at 4000 g for 30 minutes at a controlled temperature of 4°C. Centrifugation results in cells being fractionated and forming bands labelled $L_{0}-L_{4}$ in the transition zones between Percoll layers as shown in Fig. 1, with $L_{2}$ being the most concentrated layer i.e. corresponding to the peak of RBC density distribution in blood samples mohandas88densdistrib ; Ermolinskiy2020Oct ; nouaman2023effect . These different layers were then pipetted, washed, centrifuged and resuspended in a solution of 3 g/dL Dextran 70kDa to promote RBC aggregation with an interaction energy that is comparable to fibrinogen concentrations in the 4-5 g/L range neu2002 ; steffen2013quantification ; brust14 before using them separately in experiments. Due to the very low cell concentrations in layers $L_{0}$ and $L_{1}$, only layers $L_{2}$, $L_{3}$ and $L_{4}$ were used in experiments.
@@ II.2 Microfluidics
The core component of the microfluidics setup, whose principle was described in detail in a previous study and was demonstrated to be a reliable tool to characterize RBC disaggregation puthumana2024dissociation , is a hyperbolic constriction geometry (Fig. 2 a). This design has been effectively implemented in prior studies bento2018deformation ; faustino2019microfluidic to induce an extensional flow thanks to a channel width that decreases inversely with the coordinate $x$ in the flow direction ($w(x)\sim 1/x$) and a constant channel thickness $h$. Mass conservation implies that the cross-sectional average of the axial velocity $<u_{x}(x)>$ increases linearly along $x$ i.e. the average extension rate $d<u_{x}(x)>/dx$ is constant. Particles flowing along the central line therefore experience an extensional strain rate $\dot{\varepsilon_{f}}={dv_{x}}/{dx}$ where $v_{x}=u_{x}(x,0,0)$. A series of such constrictions with constant thickness $h=20$ $\mu$m, in the $z$ direction, inlet width($w_{0}=200$ $\mu$m) and outlet width ($w_{L}=20$ $\mu$m) was implemented, with decreasing lengths, $L$ (1000 $\mu$m, 500 $\mu$m, 250 $\mu$m, 125  $\mu$m), inducing decreasing extension rates $\dot{\varepsilon_{f}}$ for a given flow rate. The measurements of aggregate dissociation statistics presented here were conducted in the hyperbolic channel with $L=250$ $\mu$ m as in puthumana2024dissociation .
@@ III Results and Discussion
As previously reported puthumana2024dissociation , all disaggregation probabilities follow a characteristic sigmoid curve when increasing the extension rate, with a strong jump in the $80-200$ s-1 allowing to define a critical extension rate $\dot{\varepsilon}_{c}$ corresponding to the maximum rate of change in dissociation probabilities. To do so, experimental data in Fig. 3 is fitted with a sigmoid curve defined as:

### v3

@@ I Introduction
Red blood cell (RBC) aggregation is a critical factor influencing blood rheology and circulation baskurt2011 . Due to the presence of plasma proteins, RBCs tend to reversibly aggregate, forming mostly linear structures known as rouleaux favored by their flat biconcave shape. These linear structures can assemble into larger, gel-like networks at low shear rates and are mostly responsible for the strong shear-thinning of blood observed at low shear rates chien1970 , in conjunction with RBC deformability chien1967a ; lanotte16 .
@@
Many studies have used Percoll or Optiprep (iodixanol) density gradients to separate RBCs by density, aiming at correlating density changes, cell properties and aging. Percoll has been used for more than 40 years for density based separation of different kinds of cells without evidence of cell alteration. Recent works also specifically tested the deformability of RBCs comparing the characteristics of native cells and those that underwent Percoll processing and concluded that exposure to Percoll does not lead to alterations in RBC deformability Barshtein2024 , suggesting that it is a generally innocuous cell separation technique with few specific exceptions Wakefield1982 . These separation techniques show that RBC density distributions are relatively wide, ranging from 1.08 to 1.12 g/mL. Recent works have shown that RBC density significantly correlates with their dynamics, demonstrating a significant impact on RBC shape in flow nouaman2023effect , as well as migration and axial dispersion in capillaries losserand2019migration ; losserand2023axial .
@@
While there is no consensus on a strict equivalence between red cell density and age and studies show that there is a wide distribution of densities among cells of the same age, it is still accepted that on average density increases with cell age morrison1983does . For instance, Franco et al. franco2013 have shown that the density distribution of biotin-labeled RBCs determined at successive moments in time by density fractionation using Optiprep (Iodixanol) density gradients was significantly shifted to higher densities over the course of 120 days. A robust feature is that cell density is related to hemoglobin concentration, which tends to increase as cells age due to dehydration mohandas86scattering ; Vaysse88 . In addition, density increase was found to be correlated to loss of membrane surface area. Both phenomena lead to a decrease of static deformability (deformability index) and dynamic deformability (increase of relaxation time). Besides, density increase is also observed in RBC disorders associated with decreased cell deformability mohandas88densdistrib .
A few studies revealed that RBC aging leads to increased aggregation rampling2004influence . In one of the earliest study, Nordt 1983_Nordt showed that in autologous plasma, denser cells aggregated more than twice as much as younger, less dense cells, with intermediate-density cells displaying moderate aggregation. These results were extended to aggregation of density separated cells in 70 kDa dextran solutions by Nash et al. Nash1987 . However these early results are relatively qualitative in nature and do not provide a detailed assessment of aggregation characteristics other than variations in a global aggregation index chien1973ultrastructural . Various techniques have been proposed to quantify aggregation and characterize the aggregability of erythrocytes at the suspension scale. These include the classic erythrocyte sedimentation rate (ESR) test bedell1985 , techniques based on light scattering such as ektacytometry dacosta2016 or LORCA hardeman1994laser , and ultrasound scattering franceschini2010ultrasound , all of which are applied to dense suspensions. A limitation of these techniques is that they provide average measurements over the whole sample but fail to offer insight into the distribution of properties within a sample. Furthermore, these methods typically operate in simple shear flow, which does not fully represent conditions in the microcirculation. At the cellular scale, force measurements or controlled force application methods, such as micropipette aspiration techniques tozeren1989theoretical ; buxbaum1982quantitation , AFM steffen2013quantification , and more recently optical tweezers (OT) lee2017assessment ; Ermolinskiy2020Oct , allow direct investigation of cell-cell interactions. Recently, Ermolinskiy et al. Ermolinskiy2020Oct showed that the aggregation force measured by OT increases with the density of erythrocytes, both in autologous plasma and in 50 mg/ml Dextran 70 kDa solutions, confirming an increase of RBC aggregability with cell age. However, this type of time-consuming aggregation force measurement only allows to measure a limited number of aggregates and does not account for hydrodynamic stresses that RBCs experience in vivo, limiting the applicability of their results. Finally, flow cytometry has been proposed as a way to measure RBC aggregation Zhao_cytometry_aggregation2017 but also operates over a limited range of flow stresses that may not adequately represent in vivo flow conditions. Therefore, there remains a notable absence of quantitative techniques capable of analyzing the aging-induced variations in these interactions across a large number of aggregates in realistic flow conditions.
@@
In this work, we use this technique to quantify the variations in dissociation probabilities of small aggregates of healthy erythrocytes in extensional flow as a function of cell density after separating them using Percoll density gradients. We provide a detailed and quantitative description of RBC aggregate strength as a function of RBC density, confirming that aggregates formed from the denser fractions of red blood cells are significantly more robust than the average cell population and require significantly higher flow stresses for dissociation, and providing the dispersion of these values in a given sample. These results go beyond a simple correlation between RBC age and aggregation and add substantial value to the understanding of RBC aggregation dynamics in microcirculatory conditions. In addition, they demonstrate the relevance of the technique for improved evaluation of RBC aggregability in health and disease and provide valuable experimental data for the validation of numerical models.
@@ II.1 Blood sample collection and preparation
The separation of RBCs according to cell age was conducted by utilizing the density variations that arise during the ageing process Waugh1992Mar . To obtain density fractions of RBCs from a whole blood sample, we employed the Percoll density gradient centrifugation method outlined by Ermolinskiy et al. Ermolinskiy2020Oct . The Percoll solution (Cytiva 17-0891-01, Sigma-Aldrich, Taufkirchen, Germany) consists of Percoll, distilled water, and a 1.5M NaCl solution in different proportions adjusted to reach isotonicity with a constant measured osmolarity of $300\pm 5$ mOsm (detailed compositions are provided as Supplemental Material). This yields the five different density levels for the experiment. Five solutions, with densities of 1.085 g/mL, 1.092 g/mL, 1.101 g/mL, 1.107 g/mL, and 1.122 g/mL, were prepared and carefully placed one over the other in a 15 mL centrifugation tube (highest density solution at the bottom, lowest at the top, with each layer of volume of 2 mL). Finally, a 2 mL suspension of washed red blood cells, diluted to a hematocrit of 60% in PBS, was carefully layered onto the Percoll gradient solution. The sample was then centrifuged at 4000 g for 30 minutes at a controlled temperature of 4°C. Centrifugation results in cells being fractionated and forming bands labelled $L_{0}-L_{4}$ in the transition zones between Percoll layers as shown in Fig. 1, with $L_{2}$ being the most concentrated layer i.e. corresponding to the peak of RBC density distribution in blood samples mohandas88densdistrib ; Ermolinskiy2020Oct ; nouaman2023effect . These different layers were then pipetted, washed, centrifuged and resuspended in a solution of 3 g/dL Dextran 70kDa to promote RBC aggregation with an interaction energy that is comparable to fibrinogen concentrations in the 4-5 g/L range neu2002 ; steffen2013quantification ; brust14 before using them separately in experiments. Due to the very low cell concentrations in layers $L_{0}$ and $L_{1}$, only layers $L_{2}$, $L_{3}$ and $L_{4}$ were used in experiments. While establishing a direct relationship between RBC density and age is challenging due to the wide dispersion of densities among cells of the same age, as shown in previous studies, it is still possible to estimate age ranges for each density layer. This can be achieved using data from Franco et al. franco2013 in combination with cell lifespan models, such as those proposed in shrestha2016models , applying Bayes’ theorem to generate histograms of age versus density. However, due to the discrete nature of the original datasets, these estimates are not highly precise and should be interpreted with caution. Acknowledging this limitation, we estimate that the studied layers approximately correspond to the interquartile age ranges indicated in Table 1.
Layer Density Range (g/mL) Age Range (days) franco2013 ; shrestha2016models
$L_{2}$ (light) 1.092–1.101 $40\pm 30$
$L_{3}$ (Intermediate) 1.101–1.107 $53\pm 28$
$L_{4}$ (dense) 1.107–1.122 $58\pm 28$
Table 1: Density layers studied in the experiment and corresponding density ranges and estimated age ranges.
Note that during the density separation process, RBCs were exposed to Percoll for no more than two hours. While Percoll has been shown to promote aggregation when cells are suspended in this medium maurer2022continuous , it is completely removed after washing individual layers in PBS, with no effect on aggregability in the experiment. To rule out any potential impact of Percoll on aggregation and RBC mechanical properties, we independently verified that RBCs exposed to Percoll showed no measurable difference in aggregation compared to RBCs from the same sample that had not been in contact with Percoll (see Supplemental Material). Layers were studied in a random order and kept at 4°C while waiting to be studied in the microfluidic setup for a time course of 2-4 hours, making up a total delay between collection and study of no more than 6 hours.
@@ II.2 Microfluidics
The core component of the microfluidics setup, whose principle was described in detail in a previous study and was demonstrated to be a reliable tool to characterize RBC disaggregation puthumana2024dissociation , is a hyperbolic constriction geometry (Fig. 2 a). This design has been effectively implemented in prior studies bento2018deformation ; faustino2019microfluidic to induce an extensional flow thanks to a channel width that decreases inversely with the coordinate $x$ in the flow direction ($w(x)\sim 1/x$) and a constant channel thickness $h$. Mass conservation implies that the cross-sectional average of the axial velocity $<u_{x}(x)>$ increases linearly along $x$ i.e. the average extension rate $d<u_{x}(x)>/dx$ is constant. Particles flowing along the central line therefore experience an extensional strain rate $\dot{\varepsilon_{f}}={dv_{x}}/{dx}$ where $v_{x}=u_{x}(x,0,0)$. A series of such constrictions with constant thickness $h=20$ $\mu$m, in the $z$ direction, inlet width($w_{0}=200$ $\mu$m) and outlet width ($w_{L}=20$ $\mu$m) was implemented, with decreasing lengths, $L$ (1000 $\mu$m, 500 $\mu$m, 250 $\mu$m, 125  $\mu$m), inducing increasing extension rates $\dot{\varepsilon_{f}}$ for a given flow rate. The measurements of aggregate dissociation statistics presented here were conducted in the hyperbolic channel with $L=250$ $\mu$ m as in puthumana2024dissociation .
@@ III Results and Discussion
As previously reported puthumana2024dissociation , all disaggregation probabilities follow a characteristic sigmoid curve when increasing the extension rate. An assessment of inter-donor variability shows that differences between healthy samples are minimal, with all data points aligning on the same master curve, regardless of donor or the order in which different layers were studied. This also confirms that fluctuations in the delay between sample collection and analysis had no significant effect within the time frame of our experiments (see Supplemental Material). A strong jump observed around the extension rates of $80-200$ s-1 allows to define a critical extension rate $\dot{\varepsilon}_{c}$ corresponding to the maximum rate of change in dissociation probabilities. To do so, experimental data in Fig. 3 is fitted with a sigmoid curve defined as: